OpenStudy (anonymous):

Solve for x: 1. 42y+4 2. 2x+6 3.1.6 + 2.9x

OpenStudy (anonymous):

are they all equal?

OpenStudy (anonymous):

they are all equal to 0 meaning 42y+4=0

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

2x+6=0 2x=-6 x=-6/2 x=-3

OpenStudy (anonymous):

but wait, sorry I forgot to add something, you need to also curve sketch this: f(x)=x^3+3x^2-2

OpenStudy (anonymous):

Post it in question box

OpenStudy (anonymous):

I did.